Harris Corp. is a technology start-up and is in its second year of operations. The company didn’t purchase any assets this year but purchased the following assets in the prior year:

Placed in
Asset Service Basis
Office equipment August 14 $ 13,800
Manufacturing equipment April 15 106,000
Computer system June 1 54,000
Total $ 132,000

Harris did not know depreciation was tax deductible until it hired an accountant this year and didn’t claim any depreciation deduction in its first year of operation. (Use MACRS Table 1 and Table 2.) (Round your intermediate dollar calculations and final answer to the nearest whole dollar amount.)

a. What is the maximum amount of depreciation deduction Harris Corp. can deduct in its second year of operation?

Answer is complete but not entirely correct.

Depreciation expense   

b) What is the basis of the office equipment at the end of the second year?

Basis of office equipment   
Answer #1

a)

Since company is in second year of operation and not charge depreciation in its 1st year so they can claim depreciation of 1st year in the second accounting period as follows:

1st year Depreciation:

Depreciation on:

office equipment    13800 *15/100= 2070*5months/12=863

Manufacturing equipment 106000*15/100=15900*9months/12=11925

Computer 54000*30/100=16200*7months/12=9450

2nd year depreciation :

office eqp     13800-863=12937*15/100=1941

Manuf Eqp 106000-11925=94074*15/100=14111

Comp 54000-9450=44550*30/100=13365

Total Depreciation allowable is 863+11925+9450+1941+14111+13365=51655

b)The basis of office equipment at the end of 2nd year is 13800-863-1941=10996

By using written down value method.